OpenStudy (anonymous):

A certain radioactive isotope has a half-life of 11 days. If one is to make a table showing the half-life decay of a sample of this isotope from 32 grams to 1 gram; list the time (in days, starting with t = 0) in the first column and the mass remaining (in grams) in the second column, which type of sequence is used in the first column and which type of sequence is used in the second column?

OpenStudy (tylerd):

\[f(t)=ab^t\]

OpenStudy (tylerd):

\[f(0)=ab^0=32\] \[b^0=1\] anything to the power of 0 equals 1 solve for a we get a=32

OpenStudy (tylerd):

\[f(11)=32(b)^{11}=16\]

OpenStudy (tylerd):

\[b^{11}=\frac{ 16 }{ 32 }=\frac{ 1 }{ 2 }\]

OpenStudy (tylerd):

\[(b^{11})^{1/11}=(\frac{ 1 }{ 2 })^{1/11}=b\]

OpenStudy (anonymous):

i am so lost

OpenStudy (tylerd):

well it wants you to make a table so lets start with that then.